OpenStudy (anonymous):

Solve for x: -2 (x+3) = -2 (x+1) -4 A. 2 B. 3 C. all real numbers D. no solution

OpenStudy (anonymous):

first multiply thru to remove the negative signs 2 (x-3) = 2 (x-1) +4 then divide each side by 2 x-3 = x-1+4 move the -3 to the other side by adding 3 to each side x = x -1+4+3 ... add x = x +6 can a number equal itself plus 6? no, so your answer is?